OpenStudy (anonymous):

Can someone check my answer? For y = x2 − 4x + 3, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences to receive full credit.

OpenStudy (anonymous):

My Answer: 1)The parabola opens up because it is positive. 2)If a parabola open up, then we automatically know the vertex will be a minimum. 3)Find the vertex: x = -(-4)/2(1) =4/2 =2 y = 2^2 – 4(2) + 3 y = 4 – 8 + 3 y = -1 The vertex is (2, -1) 4)Find the x-intercepts: 0 = x^2 − 4x + 3 0 = (x -3)(x -1) x = 3 or x = 1 The x-intercepts are (3,0) and (1,0) 5)The graph shows a symmetrical parabola with a vertex minimum below the x-axis at the point of (2,-1), halfway between the x-axis intercepts at x = 1 and x = 3, with a single, vertical axis of symmetry at x = 2.

sam (.sam.):

You're good to go :)

OpenStudy (anonymous):

okay, thanks :)